OpenStudy (anonymous):

solving inequalities Will give medal 2|x+1/3|<9

OpenStudy (anonymous):

do you mean 2|(x+1)/3| < 9 or 2|x+(1/3)| < 9?

OpenStudy (anonymous):

They way I put it is the way the asked it so i have no idea

OpenStudy (anonymous):

okay, then i understand (it's the latter) first thing we have to do is isolate the absolute value expression because we can't simplify its internals that can be done by dividing by 2 on both sides 2|x + 1/3| < 9 |x + 1/3| < 9/2 now the way we deal with the absolute value is by putting it in this form: -9/2 < x + 1/3 < 9/2 we solve for x by subtracting 1/3 from everything (9/2 = 27/6, 1/3 = 2/6) -29/6 < x < 25/6 Weird numbers tho

OpenStudy (anonymous):

Still confused.

OpenStudy (gorv):

2|x+1/3|<9

OpenStudy (gorv):

divide both side by 2

OpenStudy (gorv):

|x+1/3|<9/2

OpenStudy (gorv):

now first tell me ..u know abt modulus ...what to do remove mod ??

OpenStudy (gorv):

@alexisp1